OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#11142 — FT Collect Paris
Incident Report for Web Cloud
Resolved
The LNS in Paris has lost interconnection with the FT gateway.
Sessions will be remounted on the backup LNS in Roubaix.

Update(s):

Date: 2014-07-10 08:44:04 UTC
The fibre has been fixed.
The FT collect is once again functional on the LNS in Paris.

Customer sessions will be automatically switched over to their original LNS at 04:00 (French time).


Date: 2014-07-10 08:41:28 UTC
Following a fibre disconnect on the optical loop in Paris, signal is no longer reaching the TH2 site.
The optical loop is redundant but it seems that a loop compenent did not properly manage the disconnect.
We will investigate.

Only the FT collect is impacted.

Security is doing its job for subscribers on OVH NRAs in Paris.
Posted Jul 10, 2014 - 08:34 UTC